# Neural Guided Diffusion Bridges

### v2

@@ 3.2 Guided proposal induced by linear process
We now specialize to the case where the auxiliary process is linear, i.e.
$\mathrm{d}\tilde{X}_{t}=\tilde{b}(t,\tilde{X}_{t})\mathrm{d}t+\tilde{\sigma}(t)\mathrm{d}W_{t},$ (9a)
$\tilde{b}(t,x):=\beta(t)+B(t)x.$ (9b)
Let $\tilde{A}$ denote the infinitesimal generator of the process $\tilde{X}$. Since $\tilde{h}$ solves $(\partial_{t}+\tilde{A}\tilde{)}h=0$, we can replace $(\partial_{t}+\mathcal{A})\tilde{h}$ by $(\mathcal{A}-\tilde{\mathcal{A}})\tilde{h}$. This gives
$\Psi_{t}(X)=\exp\left(\int^{t}_{0}G(s,X_{s})\mathrm{d}s\right),\quad t\leq T$ (10a)
$G(s,x):=\left\langle b(s,x)-\tilde{b}(s,x),\tilde{r}(s,x)\right\rangle$
$-\frac{1}{2}\mathrm{Tr}\left([a(s,x)-\tilde{a}(s)]\left[\tilde{H}(s)-(\tilde{r}\tilde{r}^{T})(s,x)\right]\right).$ (10b)
Here, $\tilde{H}(s)$ is the negative Hessian of $\log\tilde{h}(s,x)$, which turns out to be independent of $x$, $\tilde{a}(s)=(\tilde{\sigma}\tilde{\sigma}^{T})(s)$. Under the choice of $q(v\mid y)=\psi(v;Ly,\Sigma)$ and $\tilde{X}$, $\tilde{H}$ and $\tilde{r}$ are given by
$\tilde{H}(t)$ $=L^{T}(t)M(t)L(t)$ (11a)
$\tilde{r}(t,x)$ $=L^{T}(t)M(t)(v-u(t)-L(t)x),$ (11b)
where $M(t)=(M^{{\dagger}}(t))^{-1}$ and $L$, $M^{\dagger}$ and $u$ satisfy the backward ordinary differential equations (ODEs) (See Mider et al. [2021, Theorem 2.4]):
$\mathrm{d}L(t)$ $=-L(t)B(t)\mathrm{d}t,\quad L(T)=L$ (12a)
$\mathrm{d}M^{{\dagger}}(t)$ $=-L(t)\tilde{a}(t)L^{T}(t)\mathrm{d}t,\quad M^{{\dagger}}(T)=\Sigma$ (12b)
@@ 3.3 Choice of linear process
    In case of conditioning on the event $\{LX_{T}=v\}$ –so no extrinsic noise on the observation– the triplet needs to satisfy certain “matching conditions” to ensure $\mathbb{P}^{\star}\ll\mathbb{P}^{\circ}$. For uniformly elliptic diffusions, this only affects $\tilde{\sigma}$. In case $L=\mathbf{I}_{d}$, so the conditioning is on the full state, $\tilde{\sigma}$ should be chosen such that $\tilde{a}(T)=a(T,x_{T})$. Hence, for this setting, we can always ensure absolute continuity. For the partially observed case, it is necessary to assume that $a$ is of the form $a(t,x)=s(t,Lx)$ for some matrix values map $s$. In that case, it suffices to choose $\tilde{a}$ such that $L\tilde{a}(T)L^{T}=Ls(T,v)L^{T}$. In case the diffusivity does not depend on the state, a natural choice is to take $\tilde{\sigma}=\sigma$ to guarantee absolute continuity.
    For hypo-elliptic diffusions, the restrictions are a bit more delicate. On top of conditions on $\tilde{\sigma}$, it is also required to match certain properties in the drift by choice of $B$. In the examples that we consider in Section 5 we have ensured these properties are satisfied. We refer to Bierkens et al. [2020] for precise conditions and a host of examples.
@@ 3.4 Strategies for improving upon guided proposals
  • If we write the guided path as functional of the driving Wiener process, one can update the driving Wiener increments using the pCN within a Metropolis-Hastings algorithm. Details are provided in Appendix A.1, see also the discussion in Mider et al. [2021].
  • Devising better choices for $(\beta,B,\tilde{\sigma})$.
@@ 4 Methodology
4.1 Neural guided bridge
For a specific diffusion model, it may be hard to specify the maps $B$, $\beta$ and $\tilde{\sigma}$, which may lead to a guided proposal whose realizations look rather different from the actual conditioned paths. For this reason, we propose to adjust the dynamics of the guided proposal by adding a learnable bounded term $\sigma(t,x)\vartheta_{\theta}(t,x)$ to the drift. Specifically, let $\vartheta_{\theta}\colon[0,T]\times\mathbb{R}^{d}\to\mathbb{R}^{d}$ be a function parameterized by $\theta\in\Theta$, where $\Theta$ denotes the parameter space. Assume $x\mapsto\vartheta_{\theta}(t,x)$ is Lipschitz continuous, uniformly over $t\in[0,T]$ and that $\mathbb{P}\left(\int_{0}^{T}\|\vartheta_{\theta}(t,X_{t})\|^{2}\mathrm{d}t<\infty\right)=1$. Let
$\kappa_{T}\coloneq\exp\left(\int_{0}^{T}\vartheta^{T}_{\theta}(t,X_{t})\mathrm{d}W_{t}^{\circ}-\frac{1}{2}\int_{0}^{T}\|\vartheta_{\theta}(t,X_{t})\|^{2}\mathrm{d}t\right),$ (13)
and assume $\mathbb{E}[\kappa_{T}]=1$. We refer to Liptser and Shiryaev [2001, Chapter 6] for easily verifiable conditions on $\vartheta_{\theta}$ to ensure this assumption is met. Boundedness of $\vartheta_{\theta}$ for example suffices. Define a new probability measure $\mathbb{P}^{\bullet}$ on $(\Omega,\mathcal{F}_{T})$ by
$\mathrm{d}\mathbb{P}^{\bullet}\coloneq\kappa_{T}\mathrm{d}\mathbb{P}^{\circ}$ (14)
By Girsanov’s theorem, the process $W_{t}^{\bullet}\coloneq W^{\circ}_{t}-\int_{0}^{t}\vartheta_{\theta}(s,X_{s})\mathrm{d}s$ is a $\mathbb{P}^{\bullet}$-Wiener process. Therefore, under $\mathbb{P}^{\bullet}$, the process $X$ satisfies the SDE
$\mathrm{d}X_{t}=b^{\bullet}_{\theta}(t,X_{t})\mathrm{d}t+\sigma(t,X_{t})\mathrm{d}W^{\bullet}_{t},$
$b^{\bullet}_{\theta}(s,x)=b^{\circ}(s,x)+\sigma(s,x)\vartheta_{\theta}(s,x),\quad X_{0}=x_{0}.$ (15)
We also need Assumption 1 to ensure Equation (15) has a strong solution $X_{t}^{\bullet}$.
@@ 4.1 Neural guided bridge
There exist some constants $C,C^{\prime}>0$, such that for any $x,y\in\mathbb{R}^{d}$,
$\|\vartheta_{\theta}(t,x)-\vartheta_{\theta}(t,y)\|\leq C\|x-y\|,$ (16)
$\|\vartheta_{\theta}(t,x)\|^{2}\leq C^{\prime}(1+\|x\|^{2})$ (17)
We propose to construct $\vartheta_{\theta}$ as a learnable neural network, whose goal is to approximate the difference of drift coefficients. When $\vartheta_{\theta_{\text{opt}}}=\sigma^{T}(r-\tilde{r})$, the discrepancy between $\mathbb{P}^{\bullet}$ and $\mathbb{P}^{\star}$ vanish. In practice, when parameterization is realized as a finite neural network, the Lipschitz continuity can be achieved by plugging in enough smooth activation functions like sigmoid or tanh, and weight normalization or gradient clipping can prevent extreme growth on $x$. In the experiment part, we deploy tanh as activations and gradient clipping by the norm of 1.0 throughout all the experiments to fulfil such conditions.
However, directly approaching $\sigma^{T}(r-\tilde{r})$ is non-trivial as $r$ is still intractable. To address this, we use a variational approximation where the set of measures $\{\mathbb{P}^{\bullet}_{\theta};\,\theta\in\Theta\}$ provides a variational class for approximating $\mathbb{P}^{\star}$. The following theorem shows that minimizing $\theta\mapsto\mathrm{D_{KL}}(\mathbb{P}^{\bullet}_{\theta}||\mathbb{P}^{\star})$ is equivalent to minimizing $L(\theta)$ as defined below.
@@
If we define
$L(\theta)\coloneq\mathbb{E}^{\bullet}\int_{0}^{T}\left\{\frac{1}{2}\|\vartheta_{\theta}(t,X_{t})\|^{2}-G(t,X_{t})\right\}\mathrm{d}t,$ (18)
@@
$\mathrm{D_{KL}}{(\mathbb{P}^{\bullet}_{\theta}||\mathbb{P}^{\star})}=L(\theta)+\log\frac{\tilde{h}(0,x_{0})}{h(0,x_{0})},$ (19)
with $G$ as defined in Equation (10b).
The proof is given in Appendix A.2. Note that under $\mathbb{P}^{\bullet}$ the law of $X$ depends on $\theta$ and therefore the dependence of $L$ on $\theta$ is via both $\vartheta_{\theta}$ and the samples from $X$ under $\theta$-parameterized $\mathbb{P}^{\bullet}$. We conclude that $L$ is lower bounded by $-\log\frac{\tilde{h}(0,x_{0})}{h(0,x_{0})}$. Let $\theta_{\text{opt}}$ be a local minimizer of $L(\theta)$, then $\mathrm{D_{KL}}{(\mathbb{P}^{\bullet}_{\theta_{\text{opt}}}||\mathbb{P}^{\star})}=0$, which implies $\mathbb{P}^{\bullet}_{\theta_{\text{opt}}}=\mathbb{P}^{\star}$. Heng et al. [2022] applied a similar variational formulation, however, they used unconditional samples to estimate the KL divergence, which can be less efficient when conditioned on rare events.
In general $h(0,x_{0})$ cannot be evaluated. However, for simple examples, we can evaluate it and use it as a check to see that the trained neural network is optimal (see Section 5.1).
@@ 4.2 Reparameterization and numerical implementation
Optimizing $L(\theta)$ by gradient descent requires sampling from a parameterized distribution $\mathbb{P}^{\bullet}_{\theta}$ and backpropagating the gradients through the sampling. To estimate the gradient, We use the reparameterization trick proposed in Kingma and Welling [2022]. Specifically, the existence of a strong solution $X^{\bullet}$ to Equation (15) means that there is a measurable map $\phi_{\theta}:\mathcal{C}([0,T],\mathbb{R}^{m})\to\mathcal{C}([0,T],\mathbb{R}^{d})$, such that $X^{\bullet}=\phi_{\theta}(W^{\bullet})$, with $W^{\bullet}$ a $\mathbb{P}^{\bullet}$-Wiener process. Here, we have dropped the dependence of $\phi_{\theta}$ on the initial condition $x_{0}$ as it is fixed throughout. The objective Equation (18) can be then rewritten as:
@@
Choose a finite discrete time grid $\mathcal{T}\coloneq\{t_{m}\}_{m=0,1,\dots,M}$, with $t_{0}=0,t_{M}=T$. Let $X^{\bullet}_{t_{m}},W^{\bullet}_{t_{m}}$ be the evaluations of $X^{\bullet},W^{\bullet}$ at $t=t_{m}$ respectively, $\{x^{\bullet(n)}_{t_{m}}\},\{w^{\bullet(n)}_{t_{m}}\},n=1,\dots,N$ be collections of samples of $X^{\bullet}_{t_{m}},W^{\bullet}_{t_{m}}$, and $x^{\bullet(n)}_{t_{0}}=\phi_{\theta}(w^{\bullet(n)}_{t_{0}})=x_{0}$. Then Equation (20) can be approximated as:
@@
In practice, $x^{\bullet(n)}_{t_{m}}=\phi_{\theta}(w^{\bullet(n)}_{t_{m}})$ is implemented as a numerical SDE solver $f_{\theta}(w^{\bullet(n)}_{t_{m}},t_{m-1},x^{\bullet(n)}_{t_{m-1}})$ that takes the previous step $(t_{m-1},x^{\bullet(n)}_{t_{m-1}})$ as additional arguments. As $x^{\bullet(n)}_{t_{m-1}}$ also depends on $\theta$, the gradient with respect to $\theta$ needs to be computed recursively. Leveraging automatic differentiation frameworks, all gradients can be efficiently recorded in an acyclic computational graph during the forward integration, enabling fast and accurate backpropagation for updating $\theta$. While the complexity of backpropagation scales linearly with the number of solving steps $M$ and quadratically with the dimension of the process $d$—a property inherent to gradient-based optimization methods—our approach remains highly efficient for moderate-dimensional problems and provides a robust foundation for further scalability improvements. Li et al. [2020] proposed an alternative method to efficiently estimate the gradients called stochastic adjoint sensitivity method. However, this approach requires the SDE to be in Stratonovich form, whereas this paper focuses on Itô diffusions and their conditioning. Although Stratonovich and Itô formulations are convertible, exploring this conversion warrants separate investigation. Further details about the gradient computation can be found in Appendix A.3 and the numerical algorithm is shown as Algorithm 1
Algorithm 1 Neural guided bridge training
@@
Solve Equation (12) on $\mathcal{T}$ backwards, obtain $\{\tilde{H}(t_{m})\}$,$\{\tilde{r}(t_{m},\cdot)\}$ using Equation (11).
@@
         Solve Equation (15) on $\mathcal{T}$ with $w^{\bullet(n)}=\{w^{\bullet(n)}_{t_{m}}\}$, obtain $\{x^{\bullet(n)}_{t_{m}}\}$.
@@
     Approximate $L(\theta$) by Equation (21).

### v3

@@ 3.3 Guided Proposal Induced by Linear Process
The choice of the auxiliary process $\tilde{X}$, which determines $\tilde{h}$ and hence the guided process, offers some flexibility, as long as the conclusion of Theorem 2 applies. We now specialize to the case where the process $\tilde{X}$ solves a linear SDE, as in this case $\tilde{h}$ can be obtained by solving a finite-dimensional system of ordinary differential equations (ODEs). Specifically, we assume $\tilde{X}$ solves:
$\mathrm{d}\tilde{X}_{t}=\{\beta(t)+B(t)\tilde{X}_{t}\}\,\mathrm{d}t+\tilde{\sigma}(t)\,\mathrm{d}W_{t},\quad\tilde{X}_{0}=x_{0},$ (10)
with two $x$-independent maps $\beta:[0,T]\to\mathbb{R}^{d},B:[0,T]\to\mathbb{R}^{d\times d}$. Let $\tilde{A}$ denote the infinitesimal generator of the process $\tilde{X}$, and $\tilde{b}(s,x)\coloneq\beta(s)+B(s)x$. Since $\tilde{h}$ solves $(\partial_{t}+\tilde{A})\tilde{h}=0$, we can replace $(\partial_{t}+\mathcal{A})\tilde{h}$ by $(\mathcal{A}-\tilde{\mathcal{A}})\tilde{h}$. This gives
$\Psi_{t}(X)=\exp\left(\int^{t}_{0}G(s,X_{s})\mathrm{d}s\right),\quad t\leq T,$ (11)
$G(s,x):=\left\langle b(s,x)-\tilde{b}(s,x),\tilde{r}(s,x)\right\rangle-\frac{1}{2}\mathrm{tr}\left([a(s,x)-\tilde{a}(s)]\left[\tilde{H}(s)-(\tilde{r}\tilde{r}^{\top})(s,x)\right]\right).$ (12)
Here, $\tilde{H}(s)$ is the negative Hessian of $\log\tilde{h}(s,x)$, which turns out to be independent of $x$, $\tilde{a}(s)=(\tilde{\sigma}\tilde{\sigma}^{\top})(s)$. Under the choice of $q(v\mid y)=\psi(v;Ly,\Sigma)$ and $\tilde{X}$, $\tilde{H}$ and $\tilde{r}$ are given by
$\tilde{H}(t)=L^{\top}(t)M(t)L(t),$ (13)
$\tilde{r}(t,x)=L^{\top}(t)M(t)(v-u(t)-L(t)x),$ (14)
where $M(t)=(M^{{\dagger}}(t))^{-1}$ and $L$, $M^{\dagger}$ and $u$ satisfy the system of backward ODEs (See [Mider et al., 2021, Theorem 2.4]):
$\mathrm{d}L(t)$ $=-L(t)B(t)\,\mathrm{d}t,\quad L(T)=L,$ (15a)
$\mathrm{d}M^{{\dagger}}(t)$ $=-L(t)\tilde{a}(t)L^{\top}(t)\,\mathrm{d}t,\quad M^{{\dagger}}(T)=\Sigma,$ (15b)
@@
Remark 3.
A simple choice of $\tilde{X}$ is a scaled Brownian motion, i.e.  $\tilde{X}_{t}=\tilde{\sigma}W_{t}$. If $\tilde{\sigma}\tilde{\sigma}^{\top}$ is invertible, then
$\nabla_{x}\log\tilde{p}(T,v\mid s,x)=(\tilde{\sigma}\tilde{\sigma}^{\top})^{-1}\frac{v-x}{T-s}.$ (16)
Therefore, the guided proposal has drift $b(s,x)+\sigma(s,x)\sigma^{\top}(s,x)(\tilde{\sigma}\tilde{\sigma}^{\top})^{-1}(v-x)/(T-s)$. If $\sigma(s,x)=\tilde{\sigma}$ this reduces to the guiding term proposed in Delyon and Hu [2006].
@@ 3.4 Choice of Linear Process
    In case of conditioning on the event $\{LX_{T}=v\}$ –so no extrinsic noise on the observation– the triplet needs to satisfy certain “matching conditions” (see [Bierkens et al., 2020, Assumption 2.4]) to ensure $\mathbb{P}^{\star}\ll\mathbb{P}^{\circ}$. For uniformly elliptic diffusions, this only affects $\tilde{\sigma}$. In case $L=\mathbf{I}_{d}$, so the conditioning is on the full state, $\tilde{\sigma}$ should be chosen such that $\tilde{a}(T)=a(T,x_{T})$. Hence, for this setting, we can always ensure absolute continuity. For the partially observed case, it is necessary to assume that $a$ is of the form $a(t,x)=s(t,Lx)$ for some matrix values map $s$. In that case, it suffices to choose $\tilde{a}$ such that $L\tilde{a}(T)L^{\top}=Ls(T,v)L^{\top}$. In case the diffusivity does not depend on the state, a natural choice is to take $\tilde{\sigma}=\sigma$ to guarantee absolute continuity.
    For hypo-elliptic diffusions, the restrictions are a bit more delicate. On top of conditions on $\tilde{\sigma}$, it is also required to match certain properties in the drift by choice of $B$. With the exception of the FitzHugh-Nagumo (FHN) model studied in Section 5.3, in all examples that we consider in Section 5 we have ensured these properties are satisfied. The numerical simulation results for FHN model presented in Bierkens et al. [2020] strongly suggest that the conditions posed in that paper are actually more stringent than needed for absolute continuity. For this reason, we chose the auxiliary process just as in Bierkens et al. [2020].
@@ 3.5 Strategies for Improving upon Guided Proposals
  • If we write the guided path as functional of the driving Wiener process, one can update the driving Wiener increments using the pCN within a Metropolis-Hastings algorithm. Details are provided in Section A.1, see also the discussion in Mider et al. [2021].
  • Devising better choices for $B$, $\beta$ and $\tilde{\sigma}$.
@@ 4 Methodology
4.1 Neural Guided Diffusion Bridge
For a specific diffusion process, it may be hard to specify the maps $B$, $\beta$ and $\tilde{\sigma}$, which may lead to a guided proposal whose realizations look rather different from the actual conditioned paths. For this reason, we propose to adjust the dynamics of the guided proposal by adding a learnable bounded term $\sigma(s,x)\vartheta_{\theta}(s,x)$ to the drift. Specifically, let $\vartheta_{\theta}\colon[0,T]\times\mathbb{R}^{d}\to\mathbb{R}^{d}$ be a function parameterized by $\theta\in\Theta$, where $\Theta$ denotes the parameter space. Define:
$\varkappa_{t}\coloneq\exp\left(\int_{0}^{t}\vartheta^{\top}_{\theta}(s,X_{s})\mathrm{d}W_{s}^{\circ}-\frac{1}{2}\int_{0}^{t}\|\vartheta_{\theta}(s,X_{s})\|^{2}\mathrm{d}s\right).$ (17)
We impose the following assumption on $\vartheta_{\theta}$:
@@ 4.1 Neural Guided Diffusion Bridge
The map $\vartheta_{\theta}$ is bounded and $x\mapsto\vartheta_{\theta}(s,x)$ is Lipschitz continuous, uniformly in $s\in[0,T]$.
The Lipschitz continuity ensures $\varkappa_{t}$ is a martingale with $\mathbb{E}[\varkappa_{T}]=1$.
Define a new probability measure $\mathbb{P}^{\bullet}$ on $(\Omega,\mathcal{F}_{T})$ by
$\mathrm{d}\mathbb{P}^{\bullet}\coloneq\varkappa_{T}\mathrm{d}\mathbb{P}^{\circ}$ (18)
Then by Girsanov’s theorem, the process $W_{t}^{\bullet}\coloneq W^{\circ}_{t}-\int_{0}^{t}\vartheta_{\theta}(s,X_{s})\mathrm{d}s$ is a $\mathbb{P}^{\bullet}$-Wiener process. We now define a new diffusion process $X^{\bullet}$ under $\mathbb{P}^{\bullet}$:
Definition 2.
The neural guided diffusion bridge is a diffusion process that is defined as the strong solution to the SDE:
$\mathrm{d}X_{t}=\{b(t,X_{t})+a(t,X_{t})\tilde{r}(t,X_{t})+\sigma(t,X_{t})\vartheta_{\theta}(t,X_{t})\}\,\mathrm{d}t+\sigma(t,X_{t})\,\mathrm{d}W^{\bullet}_{t},\;X_{0}=x_{0}.$ (19)
A unique strong solution $X^{\bullet}$ to Equation 19 is guaranteed due to 1.
We propose to construct $\vartheta_{\theta}$ as a learnable neural network, whose goal is to approximate the difference of drift coefficients. When $\vartheta_{\theta}=\sigma^{\top}(r-\tilde{r})$, the discrepancy between $\mathbb{P}^{\bullet}$ and $\mathbb{P}^{\star}$ vanishes. Lipschitz continuity of the neural net can be achieved by employing sufficiently smooth activation functions and weight normalization. Gradient clipping can prevent extreme growth on $x$. In our numerical experiments, we use either tanh or LipSwish [Chen et al., 2019] activations and gradient clipping by the norm of $1.0$ to fulfill such conditions.
To learn the map $\vartheta_{\theta}$, we propose a loss function derived from a variational approximation where the set of measures $\{\mathbb{P}^{\bullet}_{\theta};\,\theta\in\Theta\}$ provides a variational class for approximating $\mathbb{P}^{\star}$. The following theorem shows that minimizing $\theta\mapsto\mathrm{D_{KL}}(\mathbb{P}^{\bullet}_{\theta}||\mathbb{P}^{\star})$ is equivalent to minimizing $L$ as defined below.
@@
If we define the loss function by
$L(\theta)\coloneq\mathbb{E}^{\bullet}\int_{0}^{T}\left\{\frac{1}{2}\|\vartheta_{\theta}(s,X_{s})\|^{2}-G(s,X_{s})\right\}\mathrm{d}s,$ (20)
@@
$\mathrm{D_{KL}}{(\mathbb{P}^{\bullet}_{\theta}||\mathbb{P}^{\star})}=L(\theta)-\log\frac{\tilde{h}(0,x_{0})}{h(0,x_{0})},$ (21)
with $G(s,x)$ as defined in Equation 12.
The proof is given in Section A.2. Note that under $\mathbb{P}^{\bullet}$, the law of $X$ depends on $\theta$ and therefore the dependence of $L$ on $\theta$ is via both $\vartheta_{\theta}$ and the samples from $X$ under $\theta$-parameterized $\mathbb{P}^{\bullet}$. If $\theta_{\rm opt}$ is a local minimizer of $L$ and $L(\theta_{\rm opt})=\log\frac{\tilde{h}(0,x_{0})}{h(0,x_{0})}$, then $\theta_{\mathrm{opt}}$ is a global minimizer. This implies $\mathrm{D_{KL}}(\mathbb{P}^{\bullet}_{\theta_{\mathrm{opt}}}||\mathbb{P}^{\star})=0$ from which we obtain $\mathbb{P}^{\bullet}_{\theta_{\mathrm{opt}}}=\mathbb{P}^{\star}$. Heng et al. [2022] applied a similar variational formulation. However, contrary to our approach, in this work the drift of the bridge proposal is learned from unconditional forward samples. Not surprisingly, this can be inefficient when conditioning on a rare event.
It can be seen that $L$ is low bounded by $\log\frac{\tilde{h}(0,x_{0})}{h(0,x_{0})}$. In general, $h$ is not known in closed form, but in some simple settings it is. In such settings, we can directly inspect the value of the lower bound and assess whether the trained neural network is optimal. In Section 5.1, we provide examples for such sanity checks.
@@ 4.2 Reparameterization and Numerical Implementation
Optimizing $L$ by gradient descent requires sampling from a parameterized distribution $\mathbb{P}^{\bullet}$ and backpropagating the gradients through the sampling. To estimate the gradient, We use the reparameterization trick proposed in Kingma and Welling [2022]. Specifically, the existence of a strong solution $X^{\bullet}$ to Equation 19 means that there is a measurable map $\phi_{\theta}:\mathcal{C}([0,T],\mathbb{R}^{d_{w}})\to\mathcal{C}([0,T],\mathbb{R}^{d})$, such that $X^{\bullet}=\phi_{\theta}(W^{\bullet})$. Here, we have dropped the dependence of $\phi_{\theta}$ on the initial condition $x_{0}$ as it is fixed throughout. The objective Equation 20 can be then rewritten as:
@@
Choose a finite discrete time grid $\mathcal{T}\coloneq\{t_{m}\}_{m=0,1,\dots,M}$, with $t_{0}=0,t_{M}=T$. Let $X^{\bullet}_{t_{m}},W^{\bullet}_{t_{m}}$ be the evaluations of $X^{\bullet},W^{\bullet}$ at $t=t_{m}$ respectively, $\{x^{\bullet(n)}_{t_{m}}\},\{w^{\bullet(n)}_{t_{m}}\},n=1,\dots,N$ be collections of samples of $X^{\bullet}_{t_{m}},W^{\bullet}_{t_{m}}$, and $x^{\bullet(n)}_{t_{0}}=\phi_{\theta}(w^{\bullet(n)}_{t_{0}})=x_{0}$. Then Equation 22 can be approximated by the Monte Carlo approximation:
@@
In practice, $x^{\bullet(n)}_{t_{m}}=\phi_{\theta}(w^{\bullet(n)}_{t_{m}})$ is implemented as a numerical SDE solver $f_{\theta}(w^{\bullet(n)}_{t_{m}},t_{m-1},x^{\bullet(n)}_{t_{m-1}})$ that takes the previous step $(t_{m-1},x^{\bullet(n)}_{t_{m-1}})$ as additional arguments. As $x^{\bullet(n)}_{t_{m-1}}$ also depends on $\theta$, the gradient with respect to $\theta$ needs to be computed recursively. Leveraging automatic differentiation frameworks, all gradients can be efficiently recorded in an acyclic computational graph during the forward integration, enabling the backpropagation for updating $\theta$. While the complexity of backpropagation scales linearly with $M$ and quadratically with $d$—a property inherent to gradient-based optimization methods—our approach remains highly efficient for moderate-dimensional problems and provides a robust foundation for further scalability improvements. Further details about the gradient computation can be found in Section A.3 and the numerical algorithm is shown as Algorithm 1.
Algorithm 1 Neural guided diffusion bridge training
@@
  Solve Equation 15 on $\mathcal{T}$ backwards, obtain and store $\{\tilde{H}(t_{m})\}$,$\{\tilde{r}(t_{m},\cdot)\}$ using Equations 13 and 14.
@@
        Solve Equation 19 on $\mathcal{T}$ with $w^{\bullet(n)}=\{w^{\bullet(n)}_{t_{m}}\}$, obtain $\{x^{\bullet(n)}_{t_{m}}\}$.
@@
     Approximate $L(\theta$) by Equation 23.